OpenStudy (anonymous):

solve the following ordinary differential equations. *dv/dt=9.8-v/5 *dp/dt=0.5p-450 *dm/dt=500-0.4m

OpenStudy (agent0smith):

By separation of variables...\[\frac{ dv }{ dt} = 9.8 - \frac{ v }{ 5 } = \frac{ 1 }{ 5 }(49 - v)\] \[\frac{ dv }{ (49 - v) } = \frac{ 1 }{ 5 }dt\] Now integrate both sides

OpenStudy (agent0smith):

The others can be solved in a similar way: \[\frac{ dp }{ dt}=0.5p-450 = 0.5(p-900)\] \[\frac{ dm }{ dt }=0.4( 1250-m)\]

OpenStudy (anonymous):

thank you!

OpenStudy (agent0smith):

No prob.